There are X Blue cars in a garage. There are Y total cars in the garage. Y and X are both positive numbers. Y is also an even number. X is less than or equal to Y. The cars are taken out of the garage one by one without replacement.

1) What is the probability that the first car is blue? (I believe the answer to this one is X/Y)

2) What is the probability that the (Y/2)th car is blue?

3) What is the probability that the final car to leave the garage is blue?
Answer #1

1) probability that the first car is blue =X/Y

2) probability that the (Y/2)th car is blue =X/Y (as probability will remain same for each car to be blue if no other information is given)

3) probability that the final car to leave the garage is blue =X/Y